# Emergence of a Landau level structure in dark optical lattices

### v2

@@ Abstract
We now consider experimental signatures of the topological character of the low-energy bands of the OFL. We first consider a low-temperature Fermi gas in a harmonic trap $U({\bm{r}})$ and subjected to the flux lattice. We show in Fig. 3a the total density variation of a Fermi sea, defined by a chemical potential $\mu_{0}$ that is set inside the energy gap separating the ground and first excited band. It can be analysed within the local density approximation, which considers the system at position ${\bm{r}}$ as locally homogeneous with a chemical potential $\mu({\bm{r}})=\mu_{0}-U({\bm{r}})$. In the central disk defined by $\mu({\bm{r}})$ higher than the ground band energy, the Fermi sea forms a band insulator, with a uniform coarse-grained density (see Fig. 3c). Outside this disk, the density drops to zero, and the system develops on the edge a chiral current consistent with the bulk-edge correspondence (see Fig. 3b) [23, 24].
Figure 4: (a) In situ density profile of a Bose-Einstein condensate in a quasi-2D geometry subjected to the optical flux lattice with $V_{0}=100\,\hbar\omega_{\text{c}}$ and to an isotropic harmonic trap. We used a trapping frequency $\omega_{\text{trap}}=0.1\,\omega_{\text{c}}$ and an interaction coupling constant $g$ such that $gN/d^{2}=100\,\hbar\omega_{\text{c}}$, where $N$ is the number of atoms per spin state. (b) Velocity distribution obtained after a free expansion. The discrete Bragg peaks do not contain information on the vortex lattice structure prior expansion. (c) Velocity distribution obtained after a short pulse of quadrupolar potential $Qxy$ mapping the condensate wavefunction to the symmetric gauge. In that case, velocity distribution reveals the vortex lattice structure.
@@
This vortex lattice shares similarities with those observed in rotating Bose gases [25, 26, 27], albeit arranged here on a square lattice imposed by the OFL geometry. In practice, vortices can be first revealed as holes in the density profile [28]. One can also consider observing them after a time-of-flight expansion in order to magnify them and reveal the associated quantized phase winding. In our flux lattice implementation, the velocity distribution cannot be used as such to reveal the vortex lattice, because it only expands over several discrete Bragg peaks (see Fig. 4b) [29]. On the other hand, it was shown in [30] that any state of the lowest Landau level in the symmetric gauge expands in time-of-flight in a self-similar manner. In order to retrieve the vortex lattice configuration in time-of-flight, the condensate wavefunction should be multiplied by the factor $\exp({\rm i}\pi xy/d^{2})$ corresponding to the gauge transform from the Landau to symmetric gauge, which can be performed by apply a quadrupolar potential $Qxy$ for a short duration prior to expansion. This is confirmed by calculating the velocity distribution of the condensate wavefunction multiplied by the complex phase factor (see Fig. 4c).
@@
Different optical setups can be envisioned for implementing the proposed OFL. Since it involves a polarization pattern sitting in the plane $xy$, a first possibility is to use light beams propagating along $z$, with a spatial structure controlled by a spatial light modulators [31]. This configuration allows fine tuning of the polarization landscape, at the expense a limited spatial resolution. Another possible scheme involves laser beams propagating solely in the $xy$ plane, which typically leads to smaller unit cells, hence larger cyclotron frequencies. We present in appendix 3 such a planar laser configuration with a triangular symmetry. This dark-state OFL gives rise to similar Landau-level-like dynamics, and one expects Bose-Einstein condensates in this structure to form a triangular vortex lattice.
We also considered the effect of imperfect realization of the atom-laser coupling geometry and the influence of a residual magnetic field on the band structure. As discussed in appendix 4, fine tuning of these parameters can be used to flatten the ground band dispersion, and topological features are robust with respect to moderate uncontrolled imperfections.
@@ Acknowledgements.
We acknowledge insightful discussions with Raphael Lopes, Jérôme Beugnon, Nathan Goldman and Zoran Hadzibabic. We also thank Gediminas Juzeliunas for remarks on an earlier version of this article. This research was funded by European Union (grant TOPODY 756722 from the European Research Council), Institut Universitaire de France, and Agence Nationale de la Recherche (ANR), projects ANR-24-CE30-7961 and ANR-24-CE47-2670-01.
Appendices
1. Dark state of a $J_{g}=1\leftrightarrow J_{e}=1$ transition
In this Letter, we consider the atomic transition sketched in figure 5, connecting the ground electronic level to an excited level, both with unit angular momentum. We indicate in the figure the values of the constant factors (derived from Clebsch-Gordan coefficients) entering in the calculation of the coupling between a Zeeman state $|g,m\rangle$ of the ground level and the Zeeman state $|e,m^{\prime}=0\rangle$ of the excited level. Since the coupling factor between the states $|g,0\rangle$ and $|e,0\rangle$ is null, an atom initially prepared in the subspace $\{|g,\pm 1\rangle,|e,0\rangle\}$ will remain in this subspace if the light field only contains left- and right-handed circular polarizations.
Suppose now that the light field is monochromatic with a frequency $\omega_{L}$ larger than the resonance frequency $\omega_{A}$. Suppose also that the one-photon couplings $\kappa_{m,0}^{(ge)}$ and $\kappa_{0,m}^{(eg)}=\left[\kappa_{m,0}^{(ge)}\right]^{*}$ between the Zeeman ground states $|g,m=\pm 1\rangle$ and the Zeeman excited state $|e,0\rangle$ is at any point $\bm{r}$ much smaller than the detuning $\Delta=\omega_{L}-\omega_{A}$. One can then restrict the internal atomic dynamics to the ground states $|g,\pm 1\rangle$ manifold by introducing the two-photon Rabi frequencies
$\kappa^{(gg)}_{m_{1},m_{2}}=\frac{\kappa^{(ge)}_{m_{1},0}\,\kappa^{(eg)}_{0,m_{2}}}{\Delta}.$ (8)
This leads to the $2\times 2$ matrix $V(\bm{r})$ used in the main text [eq.(1)].
In practice, a transition $J_{g}=1\leftrightarrow J_{e}=1$ can be isolated for different alkali atoms, such as 7Li, 39K, 41K and 87Rb. More precisely, one can use a laser close to the D1 or D2 line connecting the hyperfine levels $F_{g}=F_{e}=1$. The most favourable configuration is the D1 line of 87Rb, for which the considered transition is detuned with respect the nearest line connecting $F_{g}=1$ to $F_{e}=2$ by $\simeq 800\,$MHz. For this transition (wavelength $\lambda=795\,$nm) and the laser scheme discussed in appendix 3, we get a cyclotron frequency $\omega_{\text{c}}=2\pi\times 2.0\,$kHz.
Note that our scheme can be extended to any closed three-level system in a $\Lambda$ configuration, making it relevant for a wide range of atomic gases. Note also that the possibility for the transitions $|g,\pm 1\rangle\leftrightarrow|e\rangle$ to have different frequencies $\omega_{\pm}$ provide some flexibility in the selection of laser configurations. For example, for the $J_{g}=1\leftrightarrow J_{e}=1$ transition considered above, the degeneracy between the $\omega_{\pm}$ transitions can be lifted by a static magnetic field aligned with the $z$ axis. In this case, the amplitudes given in Eqs. (2-3) can be achieved in the plane $z=0$ by choosing (with proper phases):
  • A plane wave propagating along $z$ at frequency $\omega_{+}$ to provide the constant amplitude term in $\alpha_{+}$.
  • Two linearly polarized standing waves $\cos(kx)\,\bm{u}_{y}$ and $\cos(ky)\,\bm{u}_{x}$ also at frequency $\omega_{+}$.
  • Two linearly polarized standing waves $\sin(\bm{k}_{j}\cdot\bm{r})\,\bm{\epsilon}_{j}$ at frequency $\omega_{-}$ with $\bm{\epsilon}_{j}=(\bm{u}_{x}\pm\bm{u}_{y})/\sqrt{2}$ and $\bm{k}_{j}=\frac{k}{2}\left(\bm{u}_{x}\mp\bm{u}_{y}+\sqrt{2}\bm{u}_{z}\right)$.
This scheme can be used when non-resonant excitations, i.e., the effect of light at frequency $\omega_{\pm}$ on transition $|g,\mp 1\rangle\leftrightarrow|e\rangle$, can be neglected.
Figure 5: Level scheme considered in this letter. An atom with a $J_{g}=1\leftrightarrow J_{e}=1$ transition is irradiated by a monochromatic light field with left-handed and right-handed circular polarization components. For a proper initial state, the dynamics can be restricted to the $\Lambda$ system $\{|g,\pm 1\rangle,|e,0\rangle\}$.
2. The non-degeneracy condition of OFLs
@@
In its simplest version [3], an optical flux lattice (OFL) consists in a spatially periodic coupling between two internal states $|a\rangle$ and $|b\rangle$. This coupling can be written $\hat{V}=\epsilon_{0}\hat{1}+\epsilon_{1}\,\bm{n}\cdot\bm{\hat{\sigma}}$, where the real functions $\epsilon_{0}(\bm{r})$, $\epsilon_{1}(\bm{r})$ and the unit vector $\bm{n}(\bm{r})$ are spatially periodic, and the $\hat{\sigma}_{j}$ ($j=x,y,z$) are the Pauli matrices.
@@
At any point in space, we can diagonalize $\hat{V}$ and obtain the two eigenstates |±$|\pm\rangle_{\bm{n}}$ with the energies $\epsilon_{0}\pm\epsilon_{1}$. The non-degeneracy condition stated in [3] imposes that $\epsilon_{1}$ does not cancel at any point in the unit cell. Suppose for example that $\epsilon_{1}$ is strictly positive everywhere, and consider the Berry connection $\bm{A}(\bm{r})={\rm i}\hbar\ _{\bm{n}}\langle-|\bm{\nabla}\left(|-\rangle_{\bm{n}}\right)$ and the Berry curvature $B_{z}(\bm{r})=\left(\bm{\nabla}\times\bm{A}\right)_{z}$ that emerge for an adiabatic following of the local ground state $|-\rangle_{\bm{n}}$. By definition of an OFL, the flux of $B_{z}$ across the unit cell should be non-zero. Using Stokes theorem, this flux is given by:
$\int\hskip-8.53581pt\int B_{z}\;{{\rm d}}^{2}r=\oint_{\cal C}\bm{A}\cdot{\rm d}\bm{r}+\mbox{singular contributions,}$, (9)
@@
Let us choose a gauge, for example $|-\rangle_{\bm{n}}=(-{\rm e}^{{-\rm i}\phi}\sin(\theta/2),\cos(\theta/2))^{T}$ where the spherical angles $(\theta,\phi)$ define the orientation of $\bm{n}$. The Berry connection then reads $\bm{A}=\hbar\,\sin^{2}(\theta/2)\,\bm{\nabla}\phi$, showing that a singularity of $\bm{A}$ can appear where $\phi$ is ill-defined. This occurs at the zeroes of the complex periodic function $V_{ab}=\epsilon_{1}\,{\rm e}^{-{\rm i}\phi}\,\cos\theta$, hence at places where $\theta=0$ or $\pi$. More precisely, a point where $\theta=\pi$, i.e. $V_{aa}-\epsilon_{0}=\epsilon_{1}\cos\theta<0$ will lead to a singular contribution since $\bm{A}=\hbar\,\bm{\nabla}\phi$ at this point, whereas $\bm{A}$ will vanish at a point where $\theta=0$, i.e. $V_{aa}-\epsilon_{0}>0$, and there will be no contribution to Eq. (9) in this case.
@@
The theory of complex functions indicates that the zeroes of $V_{ab}$ come by pairs, with an opposite circulation of the phase $\phi$ around the two zeroes of the pair. If $V_{aa}-\epsilon_{0}$ keeps the same sign at the location of the two zeroes, the singular contributions to Eq. (9) of the two members of the pair thus compensate each other. To obtain a non-zero flux through the unit cell of the OFL, there must be at least one pair of zeroes of $V_{ab}$ that is "rectified", i.e., for which $V_{aa}-\epsilon_{0}$ changes sign (hence $\theta$ switches from 00 to $\pi$) between the two pair members.
@@
Consider now the coupling $\hat{V}(\bm{r})$ given in Eq. (1). The off-diagonal matrix element $V_{ab}(\bm{r})=-V_{0}\,\alpha_{+}(\bm{r})\alpha_{-}^{*}(\bm{r})$ vanishes at the zeroes of the two coefficients $\alpha_{\pm}$. If the zeroes of $\alpha_{+}$ and $\alpha_{-}$ do not overlap, then the rectification described above cannot operate. All zeroes of $\alpha_{+}$ lead to a positive sign of $V_{aa}-\epsilon_{0}=V_{0}(|\alpha_{-}|^{2}-|\alpha_{+}|^{2})/2$ and their contributions to Eq. (9) thus compensate. Similarly all zeroes of $\alpha_{-}$ lead to a negative sign of $V_{aa}-\epsilon_{0}$, hence no contribution to Eq. (9) either: one cannot reach a non-zero flux in this case. The only way to circumvent this conclusion is to allow at least some partial overlap between the zeroes of $\alpha_{+}$ and $\alpha_{-}$. However in this case, $\hat{V}$ vanishes at these common zeroes, and the two eigenstates of $\hat{V}$ coincide. As we show in the main text, this local degeneracy does not prevent the achievement of a series of topological bands.
3. Triangular version of the dark-state lattice
Figure 6: (a) Bandspectrum of a dark-state OFL with a triangular symmetry, given in Eq. (12). We recognize a series of almost equally-spaced bands at low energy, akin to Landau levels. (b) Density profile of a Bose-Einstein condensate in the OFL with triangular symmetry. We used a light shift amplitude $V_{0}=100\,\hbar\omega_{\text{c}}$, an interaction strength $gN/\lambda^{2}=5\,\hbar\omega_{\text{c}}$, and a harmonic trap frequency $\omega_{\text{trap}}=0.04\,\omega_{\text{c}}$.
We focused in the main text on a square version of the lattice. We show in this appendix that the same idea can be implemented in a triangular geometry. Consider the lattice formed by superposing three standing waves in the $xy$ plane, with an in-plane linear polarization $\bm{\epsilon}_{i}$, and a $\sigma_{+}$ polarized running wave propagating along the $z$ axis. All beams are assumed to have the same frequency $\omega$ and the resulting electric field reads, up to a multiplicative constant:
@@ 3. Triangular version of the dark-state lattice
where the wave vectors $\bm{k}_{i}$ are at 120° from each other and $\bm{\epsilon}_{i}=\hat{\bm{k}_{i}}\times\hat{\bm{z}}$. The phases $\varphi_{i}$ are chosen as $\varphi_{1}=0$, $\varphi_{2}=2\pi/3$, $\varphi_{3}=4\pi/3$ and the relative amplitude $a_{0}$ of the beam propagating along $z$ is set at the value $a_{0}=3/2$. As for the lattice considered in the main text, the field can be decomposed on the ϵ±subscriptbold-italic-$\bm{\epsilon}_{\pm}$ basis
[…]
with the amplitudes $\alpha_{\pm}$ entering in the coupling matrix $\hat{V}$ of Eq. (1):
$\begin{array}[]{rcl}\alpha_{+}&=&c_{1}+c_{2}+c_{3}+\frac{3}{2}\\ \alpha_{-}&=&c_{1}-\frac{1}{2}\left(c_{2}+c_{3}\right)\ +\ {\rm i}\frac{\sqrt{3}}{2}\left(c_{2}-c_{3}\right)\end{array}$ (12)
@@
where $c_{i}=\cos(\bm{k}_{i}\cdot\bm{r})$. The zeros of $\alpha_{-}$ are found at two sets of points. The first one corresponds to the locations where all $c_{j}=1$, the second one to the locations where all $c_{j}=-1/2$. The second set coincides with the locations of the zeroes of $\alpha_{+}$, so that the coupling matrix $\hat{V}$ vanishes at these points. We recall that this local degeneracy is required for a dark-state based OFL (see appendix 2).
@@
We show in Fig. 6a the energy spectrum as a function of the light coupling strength $V_{0}$. For large enough $V_{0}$, we find a series of almost equally-spaced energy bands, with a spacing $\hbar\omega_{\text{c}}=\sqrt{3}\hbar^{2}k^{2}/(2\pi m)$ consistent with the cyclotron frequency expected for the magnetic flux density. For $V_{0}=100\,\hbar\omega_{\text{c}}$, we find a topological lowest band with a flatness ratio $\simeq 5$. An example of Bose-Einstein condensate with a triangular vortex lattice in this configuration is shown in Fig. 6b.
4. Topological robustness of the ground band
Figure 7: (a) Evolution of the flatness ratio of the lowest band (ratio between the gap and the band width) for $V_{0}=100\,\hbar\omega_{\text{c}}$ as a function of $\beta$ and $\gamma$, for $h_{z}=0$. (b) Evolution of the flatness ratio with $h_{z}$, for $\beta=\gamma=1$. The red crosses correspond to the values $\beta=\gamma=1$ and $h_{z}=0$ considered in the main text.
@@ 4. Topological robustness of the ground band
The lattice considered in the main text, described by Eqs. (1-3), depends only on the parameter $V_{0}$. Moreover the low-energy band structure is essentially independent of $V_{0}$ for large values of this parameter (figure 1). We consider here the sensitivity of our results to different kinds of imperfections, namely a modification of the light field profile and an external magnetic field. More specifically, we consider light field amplitudes $\alpha_{+}$ and $\alpha_{-}$ given by
$\alpha_{+}(\bm{r})$ $=$ $\gamma+\frac{1}{2}\left[\cos(kx)+\cos(ky)\right],$ (13)
$\alpha_{-}(\bm{r})$ $=$ $\beta\left\{\sin\left[\frac{k}{2}(x-y)\right]-{\rm i}\sin\left[\frac{k}{2}(x+y)\right]\right\}.$ (14)
@@
parametrized by $\beta$ and $\gamma$. We also study the influence of a magnetic field along $z$ giving rise to a Zeeman splitting $h_{z}m_{z}$. The band structure presented in the main text correspond to the values $\beta=\gamma=1$ and $h_{z}=0$.
We focus here on the flatness ratio of the lowest band, i.e., the ratio of the gap to the first excited band and the width of this lowest band. We plot the result in Fig. 7 for $V_{0}=100\,\hbar\omega_{\text{c}}$. We find a non-zero flatness ratio for most of the range explored in the figure ($1.50.5<\beta,\gamma<$ 1.5), indicating the robustness of the topology of the lattice. Similarly, the topological character of the band is preserved for Zeeman fields $|h_{z}|\lesssim 2\,\hbar\omega_{\text{c}}$. We also note that the band width can be significantly decreased with a fine tuning of the values of $\beta$ and $\gamma$. For $\beta=0.92$, $\gamma=0.88$ and $h_{z}=-0.25\,\hbar\omega_{\text{c}}$, this width is only $0.06\,\hbar\omega_{\text{c}}$ (flatness ratio of 13.9) instead of $0.11\,\hbar\omega_{\text{c}}$ for $\beta=\gamma=1$ and $h_{z}=0$ (flatness ratio of 7.2). This fine tuning should ease the reach of fractional quantum Hall states, for which the interaction energy dominates over the kinetic energy (i.e. the band width).
We mention that our dark-state OFL is also intrinsically robust to polarization defects, given the fact that a dark internal state can be always be defined for any monochromatic electric field at the vicinity of a $J_{g}=1$ to $J_{e}=1$ optical transition [32].

### v3

@@ Abstract
We now consider experimental signatures of the topological character of the low-energy bands of the OFL. We first consider a low-temperature Fermi gas in a harmonic trap $U({\bm{r}})$ and subjected to the OFL. We show in Fig. 3a its total density variation, assuming a chemical potential $\mu_{0}$ set in the gap separating the ground and first excited bands. In the local density approximation framework, we consider the system at position $\bm{r}$} as locally homogeneous with a chemical potential $\mu({\bm{r}})=\mu_{0}-U({\bm{r}})$. In the central disk defined by $\mu({\bm{r}})$ higher than the ground band energy, the Fermi sea forms a band insulator, with a uniform coarse-grained density (see Fig. 3c). Outside this disk, the density drops to zero, and the system develops on the edge a chiral current consistent with the bulk-edge correspondence (see Fig. 3b) [28, 29].
Figure 4: (a) In situ density profile of a Bose-Einstein condensate in a quasi-2D geometry subjected to the optical flux lattice with $V_{0}=50\,\hbar\omega_{\text{c}}$ and to an isotropic harmonic trap. We used a trapping frequency $\omega_{\text{trap}}=0.06\,\omega_{\text{c}}$ and an interaction coupling constant $g$ such that $gN/d^{2}=100\,\hbar\omega_{\text{c}}$, where $N$ is the number of atoms per spin state. (b) Velocity distribution obtained after a free expansion. The discrete Bragg peaks do not contain information on the vortex lattice structure prior expansion. (c) Velocity distribution obtained after a short pulse of quadrupolar potential QxyQxy mapping the condensate wavefunction to the symmetric gauge. In that case, velocity distribution reveals the vortex lattice structure.
@@
This vortex lattice shares similarities with those observed in rotating Bose gases [30, 31, 32], albeit arranged here on a square lattice imposed by the OFL geometry. In practice, vortices can be revealed as holes in the density profile [33]. One can also consider observing them after a time-of-flight expansion in order to magnify them and reveal the associated quantized phase winding. In our OFL implementation, the velocity distribution cannot be used as such to reveal the vortex lattice, because it only expands over discrete Bragg peaks (see Fig. 4b) [34]. On the other hand, it was shown in [35] that any state of the lowest Landau level in the symmetric gauge expands in time-of-flight in a self-similar manner. In order to retrieve the vortex lattice configuration in time-of-flight, the condensate wavefunction should be multiplied by the factor $\exp({\rm i}\pi xy/d^{2})$ corresponding to the gauge transform from the Landau to symmetric gauge, which can be performed by applying a quadrupolar potential $Qxy$ for a short duration prior to expansion. This is confirmed by calculating the velocity distribution of the condensate wavefunction multiplied by the complex phase factor (see Fig. 4c). Note that triangular vortex arrays can also be achieved with dark state lattices [19, 23].
@@
[…]
We also considered the effect of imperfect realization of the atom-laser coupling and the influence of a residual deviation from the Raman resonance transition on the band structure. As discussed in [19], fine tuning of these parameters can be used to flatten the ground band dispersion, and topological features are robust with respect to moderate uncontrolled imperfections.
@@ Acknowledgements.
We acknowledge insightful discussions with Jérôme Beugnon, Nigel Cooper, Nathan Goldman, Zoran Hadzibabic and Raphael Lopes. We also thank Gediminas Juzeliunas for remarks on an earlier version of this article. This research was funded by European Union (grant TOPODY 756722 from the European Research Council), Institut Universitaire de France, and Agence Nationale de la Recherche (ANR), projects ANR-24-CE30-7961 and ANR-24-CE47-2670-01.
Supplemental Material
Emergence of a Landau level structure
in dark optical lattices
S1. The non-degeneracy condition of OFLs
@@
In its simplest version [3], an optical flux lattice (OFL) consists in a spatially periodic coupling between two internal states $|a\rangle$ and $|b\rangle$ (not necessarily the two states $|g_{+}\rangle$ and $|g_{-}\rangle$ introduced in the main text). This coupling can be written $\hat{V}=\epsilon_{0}\hat{1}+\epsilon_{1}\,\bm{n}\cdot\bm{\hat{\sigma}}$, where the real functions $\epsilon_{0}(\bm{r})$, $\epsilon_{1}(\bm{r})$ and the unit vector $\bm{n}(\bm{r})$ are spatially periodic, and the $\hat{\sigma}_{j}$ ($j=x,y,z$) are the Pauli matrices.
@@
At any point in space, we can diagonalize $\hat{V}$ and obtain the two eigenstates $|\pm\rangle_{\bm{n}}$ with the energies $\epsilon_{0}\pm\epsilon_{1}$. The non-degeneracy condition stated in [3] imposes that $\epsilon_{1}$ does not cancel at any point in the unit cell. Suppose for example that $\epsilon_{1}$ is strictly positive everywhere, and consider the Berry connection $\bm{A}(\bm{r})={\rm i}\hbar\ _{\bm{n}}\langle-|\bm{\nabla}\left(|-\rangle_{\bm{n}}\right)$ and the Berry curvature $B_{z}(\bm{r})=\left(\bm{\nabla}\times\bm{A}\right)_{z}$ that emerge for an adiabatic following of the local ground state $|-\rangle_{\bm{n}}$. By definition of an OFL, the flux of $B_{z}$ across the unit cell should be non-zero. Using Stokes theorem, this flux is given by:
$\int\thinspace\int B_{z}\;{{\rm d}}^{2}r=\oint_{\cal C}\bm{A}\cdot{\rm d}\bm{r}+\mbox{singular contributions,}$ (S1)
@@
Let us choose a gauge, for example $|-\rangle_{\bm{n}}=(-{\rm e}^{{-\rm i}\phi}\sin(\theta/2),\cos(\theta/2))^{T}$ where the spherical angles $(\theta,\phi)$ define the orientation of $\bm{n}$. The Berry connection then reads $\bm{A}=\hbar\,\sin^{2}(\theta/2)\,\bm{\nabla}\phi$, showing that a singularity of $\bm{A}$ can appear where $\phi$ is ill-defined. This occurs at the zeroes of the complex periodic function $V_{ab}=\epsilon_{1}\,{\rm e}^{-{\rm i}\phi}\,\cos\theta$, hence at places where $\theta=0$ or $\pi$. More precisely, a point where $\theta=\pi$, i.e. $V_{aa}-\epsilon_{0}=\epsilon_{1}\cos\theta<0$ will lead to a singular contribution since $\bm{A}=\hbar\,\bm{\nabla}\phi$ at this point, whereas $\bm{A}$ will vanish at a point where $\theta=0$, i.e. $V_{aa}-\epsilon_{0}>0$, and there will be no contribution to Eq. (S1) in this case.
@@
The theory of complex functions indicates that the zeroes of $V_{ab}$ come by pairs, with an opposite circulation of the phase $\phi$ around the two zeroes of the pair. If $V_{aa}-\epsilon_{0}$ keeps the same sign at the location of the two zeroes, the singular contributions to Eq. (S1) of the two members of the pair thus compensate each other. To obtain a non-zero flux through the unit cell of the OFL, there must be at least one pair of zeroes of $V_{ab}$ that is "rectified", i.e., for which $V_{aa}-\epsilon_{0}$ changes sign (hence $\theta$ switches from 0 to $\pi$) between the two pair members.
@@
Consider now the coupling $\hat{V}(\bm{r})$ given in Eq. (1). The off-diagonal matrix element $V_{ab}(\bm{r})=V_{0}\,\alpha_{+}^{*}(\bm{r})\alpha_{-}(\bm{r})$ vanishes at the zeroes of the two coefficients $\alpha_{\pm}$. If the zeroes of $\alpha_{+}$ and $\alpha_{-}$ do not overlap, then the rectification described above cannot operate. All zeroes of $\alpha_{+}$ lead to a positive sign of $V_{aa}-\epsilon_{0}=V_{0}(|\alpha_{-}|^{2}-|\alpha_{+}|^{2})/2$ and their contributions to Eq. (S1) thus compensate. Similarly all zeroes of $\alpha_{-}$ lead to a negative sign of $V_{aa}-\epsilon_{0}$, hence no contribution to Eq. (S1) either: one cannot reach a non-zero flux in this case. The only way to circumvent this conclusion is to allow at least some partial overlap between the zeroes of $\alpha_{+}$ and $\alpha_{-}$. However in this case, $\hat{V}$ vanishes at these common zeroes, and the two eigenstates of $\hat{V}$ coincide. As we show in the main text, this local degeneracy does not prevent the achievement of a series of topological bands.
S2. Ideal Chern band character
Recent theoretical developements introduced a class of ‘ideal’ Chern bands, which share the same algebraic structure as Landau levels despite a non-uniform Berry curvature [40, 36, 27, 41]. These bands are characterized by a specific structure of the quantum geometrical tensor $\mathcal{Q}_{n,{\bm{q}}}]$, defined by
$\mathcal{Q}_{n,{\bm{q}}}]_{ij}=\langle\partial_{q_{i}}\psi_{n,{\bm{q}}}\rvert(1-\lvert\psi_{n,{\bm{q}}}\rangle\langle\psi_{n,{\bm{q}}}\rvert)\lvert\partial_{q_{j}}\psi_{n,{\bm{q}}}\rangle,$
where $\lvert\psi_{n,{\bm{q}}}\rangle$ is the Bloch state of band $n$ and quasi-momentum $\bm{q}$}.
In the case of Landau levels, each band is characterized by a uniform quantum geometrical tensor given by the matrix form
$\mathcal{Q}_{n,{\bm{q}}}]=\frac{\ell^{2}}{2}\begin{pmatrix}1&-{\rm i}\\ {\rm i}&1\end{pmatrix}$,
where $\ell$ is the magnetic length. It satisfies the identity $\mathcal{Q}_{n,{\bm{q}}}]\bm{\epsilon}_{-}=\bm{0}$ (where $\bm{\epsilon}_{\pm}=({\bm{e}}_{x}\pm{\rm i}\,{\bm{e}}_{y})/\sqrt{2}$), which physically means that the dynamics of the momentum component $q_{x}-{\rm i}q_{y}$ is suppressed. This property is at the heart of the possibility to write Landau levels wavefunctions in terms of holomorphic functions.
More generally, ideal Chern bands are characterized by the existence of a constant null vector $\bm{w}$}, i.e. $\mathcal{Q}_{n,{\bm{q}}}]{\bm{w}}=\bm{0}$ for all $\bm{q}$}, which provides the same algebraic structure than Landau levels [36].
@@
S3. Triangular version of the dark-state lattice
Figure S1: (a) Bandspectrum of a dark-state OFL with a triangular symmetry, given in Eq. (S2). We recognize a series of almost equally-spaced bands at low energy, akin to Landau levels. (b) Density profile of a Bose-Einstein condensate in the OFL with triangular symmetry. We used a light shift amplitude $V_{0}=100\,\hbar\omega_{\text{c}}$, an interaction strength $gN/\lambda^{2}=5\,\hbar\omega_{\text{c}}$, and a harmonic trap frequency $\omega_{\text{trap}}=0.04\,\omega_{\text{c}}$.
We focused in the main text on a square version of the lattice. We show in this appendix that the same idea can be implemented in a triangular geometry. Consider the lattice formed by superposing three standing waves in the $xy$ plane, superimposed with a running wave propagating along the $z$ axis, so that the amplitudes $\alpha_{\pm}$ entering in the coupling matrix $\hat{V}$ of Eq. (1) read:
$\begin{array}[]{rcl}\alpha_{+}&=&c_{1}+c_{2}+c_{3}+\frac{3}{2}\\ \alpha_{-}&=&c_{1}-\frac{1}{2}\left(c_{2}+c_{3}\right)\ +\ {\rm i}\frac{\sqrt{3}}{2}\left(c_{2}-c_{3}\right)\end{array}$ (S2)
@@
where $c_{i}=\cos(\bm{k}_{i}\cdot\bm{r})$ and the wave vectors $\bm{k}_{i}$ are at 120° from each other. The zeros of $\alpha_{-}$ are found at two sets of points. The first one corresponds to the locations where all $c_{j}=1$, the second one to the locations where all $c_{j}=-1/2$. The second set coincides with the locations of the zeroes of $\alpha_{+}$, so that the coupling matrix $\hat{V}$ vanishes at these points. We recall that this local degeneracy is required for a dark-state based OFL (see appendix S1).
@@
We show in Fig. S1a the energy spectrum as a function of the light coupling strength $V_{0}$. For large enough $V_{0}$, we find a series of almost equally-spaced energy bands, with a spacing $\hbar\omega_{\text{c}}=\sqrt{3}\hbar^{2}k^{2}/(2\pi m)$ consistent with the cyclotron frequency expected for the magnetic flux density. For $V_{0}=100\,\hbar\omega_{\text{c}}$, we find a topological lowest band with a flatness ratio $\simeq 5$. An example of Bose-Einstein condensate with a triangular vortex lattice in this configuration is shown in Fig. S1b.
S4. Topological robustness of the ground band
Figure S2: (a) Phase diagram for the ground energy band of the OFL as a function of the parameters $\beta$ and $h_{z}$ for $V_{0}=100\,\hbar\omega_{\text{c}}$. The phase (I) exhibits a gapped ground band with a Chern number $\mathcal{C}=1$. The phases (II) and (III) are both gapped with a Chern number $\mathcal{C}=0$. In the phase (IV), the ground band overlaps with excited Bloch bands. The blue lines indicate the gapless transition lines separating the different phases. (b) Evolution of the flatness ratio with $\beta$ and $\delta$. The red cross corresponds to the values $\beta=1$ and $\delta=0$ considered in the main text. A maximum flatness ratio of about 130 is obtained for $\beta=0.65$ and $\delta=2.9\,\omega_{\text{c}}$ (barely visible on the plot due the strong sensitivity of high flatness ratios to $\beta$ and $\delta$).
@@
The lattice considered in the main text, described by Eqs. (1-3), depends only on the parameter $V_{0}$. Moreover the low-energy band structure is essentially independent of $V_{0}$ for large values of this parameter (see Fig. 1). We consider here the sensitivity of our results to different kinds of imperfections, namely a modification of the light field profile and an external magnetic field. More specifically, we consider light field amplitudes $\alpha_{+}$ and $\alpha_{-}$ given by
$\alpha_{+}(\bm{r})$ $=$ $\sin X+{\rm i}\sin Y,$ (S3)
$\alpha_{-}(\bm{r})$ $=$ $\beta\left(\cos X+\cos Y\right),$ (S4)
@@
parametrized by $\beta$. We also study the influence of a magnetic field along $z$ giving rise to a Zeeman splitting $\hbar\delta\hat{\sigma}_{z}/2$. Equivalently, such a splitting also arises from a detuning $\delta$ with respect to the Raman resonance. The band structure presented in the main text correspond to the values $\beta=1$ and $\delta=0$.
We first consider the phase diagram of the system, more specifically the topological nature of the ground band. As shown in Fig. S2a, we find a gapped topological band with Chern number $\mathcal{C}=1$ for a wide range of parameters around $(\beta,\delta)=(1,0)$ (phase I). When increasing the magnitude of the Zeeman field, one finds transition lines at which the gap to the first excited band vanishes, marking the transition to non-topological bands (phases II and III). For weak values of $\beta$, the system becomes metallic (phase IV).
We then consider the flatness ratio of the lowest band, i.e., the ratio of the gap to the first excited band and the width of this lowest band (see Fig. S2b). Fine tuning of the parameters $\beta$ and $\delta$ allows to flatten the ground band significantly, with flatness ratios exceeding 100. This fine tuning should ease the reach of fractional quantum Hall states, for which the interaction energy dominates over the kinetic energy (i.e. the band width).